OpenStudy (anonymous):

In the abbreviation of the trigonometric form of a complex number, r times c i s of theta., what does the s represent? A. secant B. sine C. square root D. sum

OpenStudy (anonymous):

sine

OpenStudy (tkhunny):

\(cis(\theta) = \cos(\theta) + i\cdot\sin(\theta)\) "c": cosine "i": i "s": sine